OpenStudy (anonymous):

what number must you add to complete the square? x^2+4x=3 A. 8 B. 4 C. 2 D. 16

OpenStudy (alexandervonhumboldt2):

x^2+4x=3 x^2+4x-3=0 (a+b)^2=a^2+2ab+b^2 x^2+2*2*x-3=0 x^2+2*2*x+2^2-3-2^2=0 x^2+4x+4-3-4=0 (x+2)^2-7=0 we add 4 to complete the square

OpenStudy (anonymous):

thanks